Jeremy inherits $18,000 and invests part of the money in bonds with an intrest rate of 11%. The remainder of the money is in savinga at a 9% rate. What amount has he invested at each rate if he receives $1,820 in interest for 1 year?
Answer by checkley71(8403)

.11X+.09(18000-X)=1820
.11X+1620-.09X=1820
.02X=1820-1620
.02X=200
X=200/.02
X=10000 INVESTED @ 11%
18000-10000=8000 INVESTED @9%
PROOF
.11*10000+.09*8000=1820
1100+720=1820
1820=1820